FOURIER TRANSFORM MICROWAVE SPECTRA OF N-METHYLFORMAMIDE

\underline{YOSHIYUKI KAWASHIMA}, TSUYOSHI USAMI, Department of Applied Chemistry, Kanagawa Institute of Technology, Atsugi, Kanagawa 243-0292, JAPAN; RICHARD D. SUENRAM, National Institute of Standards and Technology, Gaithersburg, MD 20899; and EIZI HIROTA, The Graduate University for Advanced Studies, Hayama, Kanagawa 240-0193, JAPAN.

In order to understand the dynamical behavior of the peptide bond in “peptide molecules: X-CONH-Y”, we have chosen N-methylformamide (N-MFA) as a simplest system with X = H and Y = CH_{3}. The microwave spectrum of the most stable trans form of N-MFA was studied by Fantoni and Caminati, ^{a} who assigned and analyzed only the A state of the CH_{3} internal rotation and thus reported an estimated potential barrier V_{3}. Kitano and Kuchitsu ^{b} determined the molecular structure of trans N-MFA by using electron diffraction, but they could not detect the higher energy cis conformer. We report here the rotational spectrum of both the trans and cis forms of N-MFA using FTMW spectrometers at NIST and KAIT, in the frequency region from 5 to 36 GHz, with either Ne or Ar as a buffer gas in a reservoir nozzle maintained at about 50^{o}C. We confirmed the assignment of Ref.a for the trans A state and extended the observation to low-J transitions. The spectra of four trans isotopomers ^{13}C, ^{15}N, and ^{18}O were observed to derive an r_{s} structure, which is in good agreement with the electron diffraction and ab initio results. The V_{3} value of trans is approximately 60 cm^{−1}, which is 4.5 times smaller than the value of cis form which was determined to be 279 cm^{−1}.

aA. C. Fantoni and W. Caminati J. Chem. Soc. Faraday Trans. \underline{92}, 342 (1995).

bM, Kitano and K. Kuchitsu Bull. Chem. Soc. Jpn. \underline{47}, 631 (1974).